A new, larger size piping (1/4") heat exchanger was designed for
the helium target. It was installed in September. The total cooling power
for the new helium target reached to about 500 Watts.
In January, Helium loop was reconfigured to
use two heat exchangers (loop1 and loop 3) in parallel. The total
cooling power was doubled (to 440 Watts) after the reconfiguration.
The helium loop was installed in Nov., 1999. A 3He
gas handling system was also installed. Both 3, 4He targets
were cooled down succesfully to 6.3 K and about 10 atm. It was found that
the small size (1/8") of the heat exchanger piping severely limited the
coolant flow. So the total cooling power was only about 220 Watts.
Both Hydrogen and Deuterium loops were successfully cooled down on
Sept 29, 1997.
Hydrogen loop was filled with LH2 at 19 K, 26 PSIA
Deuterium loop was filled with LD2 at 22 K, 22 PSIA
Both targets have been in stable operation condition since cooled down.
The cryotargets commissioned with beam from 10/2 to 10/8/97.
Both 15 cm LH2 and LD2 took up to 120 MicroAmps
Beam power over 700 Watts. Luminosity is 5x1038 cm-2*sec-1